# From Light-Cone to Supersonic Propagation of Correlations by Competing Short- and Long-Range Couplings

### v2

@@ Abstract
The local processes consist in the repulsive on-site interactions of strength $U$ and atomic tunneling between neighboring sites $\langle j,j^{\prime}\rangle$ with the amplitude $J$. The model provides a very good description of the dynamics realized in experiments Klinder et al. (2015); Landig et al. (2016); Hruby et al. (2018), where the atomic transition is far detuned and spontaneous decay can be neglected. The effective light-matter coupling strength $\Omega$ is controllable by a pump laser and $\delta$ is the cavity-pump detuning. The period of the lattice is twice the period of the cavity mode coupling the cavity to the density imbalance $\hat{\Delta}=\sum_{j\in A}\hat{n}_{j}-\sum_{j\in B}\hat{n}_{j}$ of a bipartite lattice with sublattices AA and $B$ Habibian et al. (2013); Niederle et al. (2016); Landig et al. (2016). Theoretical efforts characterized mostly the nature of steady states in such models Niedenzu et al. (2010); Silver et al. (2010); Fernández-Vidal et al. (2010); Li et al. (2013); Bakhtiari et al. (2015); Flottat et al. (2017); Lin et al. (2019); Himbert et al. (2019); Halati et al. (2020a); Bezvershenko et al. (2021); Sharma et al. (2022); Chanda et al. (2022), and certain aspects of their dynamics Rodríguez Chiacchio and Nunnenkamp (2018); Halati et al. (2022, 2025).
@@ A Time-dependent matrix product states method for long-range quantum systems
In order to ensure the convergence of the density-density correlations, $\mathcal{C}_{nn}$, [see main text Eq. (3) for their definition], up to an error of approximately $10^{-5}$ for the times considered in the results presented in the absence of dissipation, we chose the following convergence parameters: a maximal bond dimension of 400400 states, which ensured a truncation error of at most $5\times 10^{-8}$ at the final time, a time-step of $dtJ/\hbar=2.5\times 10^{-3}$, the local Hilbert space of the bosonic atoms is $N_{\text{bos}}=4$, and the adaptive cutoff of the local Hilbert space of the photonic mode ranged between $N_{\text{pho}}=70$ and $N_{\text{pho}}=8$. For the parameters used in the presence of dissipation a truncation error of $10^{-8}$ at the final time was achieved for the mentioned convergence parameters, the presented results are averaged over at least 750 quantum trajectories.
The numerical results for the time-evolution of the atomic model with global range interactions, Eq. (4) in the main text, were obtained with an implementation based on matrix product states of the two-site version of the time-dependent variational principle approach (TDVP) Haegeman et al. (2011, 2016), within the ITensor Library Fishman et al. (2022). The convergence was assured by the following parameters of the method: a maximal bond dimension of 250250 states, which ensured a truncation error of at most $10^{-7}$ at the final time, a time-step of $dtJ/\hbar=5\times 10^{-3}$, and the local Hilbert space of the bosonic atoms are $N_{\text{bos}}=4$.
@@ C Treating the cavity mode as a classical field
where $\xi(t)$ is a random complex number of magnitude 11 sampled from a uniform distribution at each point in time. For the initial state $\langle a\rangle(0)$ we use a uniformly sampled random complex number. We simulate different realizations of the time-dependent noise term, similar to the quantum trajectories approach, and average over the values of the computed observables for the different realizations. We obtain that in the presence of the stochastic noise, Fig. C1(e) and Fig. C1(f) it exhibits for larger distances features of the distance-independent rise of the correlations, as seen in Fig. C1(f) and in the log-scale depiction of the correlations in Fig. C2. This can be understood by the fact that the noise term appearing in Eq. (C.3) translates to a correlated noise term in the evolution of the atoms acting on all sites. However, the contribution from the distance-independent rise of the correlations is much smaller for the shown parameters compared to the full quantum dynamics. This implies that the noise term is insufficient to capture the full dynamics which requires, in particular, also the cavity-mediated transport of density fluctuations.
@@ 2 Technical details on the simulations
where $j_{0}=(L+1)/2$ is the center and the division by 22 comes from the fact that there are two sites with the same distance in 1D. The result of this calculation is shown in Fig. 3(c) in the main text.

### v3

@@ Abstract
The local processes consist in the repulsive on-site interactions of strength $U$ and atomic tunneling between neighboring sites $\langle j,j^{\prime}\rangle$ with the amplitude $J$. The model provides a very good description of the dynamics realized in experiments Klinder et al. (2015); Landig et al. (2016); Hruby et al. (2018), where the atomic transition is far detuned and spontaneous decay can be neglected. The effective light-matter coupling strength $\Omega$ is controllable by a pump laser and $\delta$ is the cavity-pump detuning. The period of the lattice is twice the period of the cavity mode coupling the cavity to the density imbalance $\hat{\Delta}=\sum_{j\in A}\hat{n}_{j}-\sum_{j\in B}\hat{n}_{j}$ of a bipartite lattice with sublattices $A$ and $B$ Habibian et al. (2013); Niederle et al. (2016); Landig et al. (2016). Theoretical efforts characterized mostly the nature of steady states in such models Niedenzu et al. (2010); Silver et al. (2010); Fernández-Vidal et al. (2010); Li et al. (2013); Bakhtiari et al. (2015); Flottat et al. (2017); Lin et al. (2019); Himbert et al. (2019); Halati et al. (2020a); Bezvershenko et al. (2021); Sharma et al. (2022); Chanda et al. (2022), and certain aspects of their dynamics Rodríguez Chiacchio and Nunnenkamp (2018); Halati et al. (2022, 2025).
@@ A Time-dependent matrix product states method for long-range quantum systems
In order to ensure the convergence of the density-density correlations, $\mathcal{C}_{nn}$, [see main text Eq. (3) for their definition], up to an error of approximately $10^{-5}$ for the times considered in the results presented in the absence of dissipation, we chose the following convergence parameters: a maximal bond dimension of $400$ states, which ensured a truncation error of at most $5\times 10^{-8}$ at the final time, a time-step of $dtJ/\hbar=2.5\times 10^{-3}$, the local Hilbert space of the bosonic atoms is $N_{\text{bos}}=4$, and the adaptive cutoff of the local Hilbert space of the photonic mode ranged between $N_{\text{pho}}=70$ and $N_{\text{pho}}=8$. For the parameters used in the presence of dissipation a truncation error of $10^{-8}$ at the final time was achieved for the mentioned convergence parameters, the presented results are averaged over at least 750 quantum trajectories.
The numerical results for the time-evolution of the atomic model with global range interactions, Eq. (4) in the main text, were obtained with an implementation based on matrix product states of the two-site version of the time-dependent variational principle approach (TDVP) Haegeman et al. (2011, 2016), within the ITensor Library Fishman et al. (2022). The convergence was assured by the following parameters of the method: a maximal bond dimension of $250$ states, which ensured a truncation error of at most $10^{-7}$ at the final time, a time-step of $dtJ/\hbar=5\times 10^{-3}$, and the local Hilbert space of the bosonic atoms are $N_{\text{bos}}=4$.
@@ C Treating the cavity mode as a classical field
where $\xi(t)$ is a random complex number of magnitude $1$ sampled from a uniform distribution at each point in time. For the initial state $\langle a\rangle(0)$ we use a uniformly sampled random complex number. We simulate different realizations of the time-dependent noise term, similar to the quantum trajectories approach, and average over the values of the computed observables for the different realizations. We obtain that in the presence of the stochastic noise, Fig. C1(e) and Fig. C1(f) it exhibits for larger distances features of the distance-independent rise of the correlations, as seen in Fig. C1(f) and in the log-scale depiction of the correlations in Fig. C2. This can be understood by the fact that the noise term appearing in Eq. (C.3) translates to a correlated noise term in the evolution of the atoms acting on all sites. However, the contribution from the distance-independent rise of the correlations is much smaller for the shown parameters compared to the full quantum dynamics. This implies that the noise term is insufficient to capture the full dynamics which requires, in particular, also the cavity-mediated transport of density fluctuations.
@@ 2 Technical details on the simulations
where $j_{0}=(L+1)/2$ is the center and the division by $2$ comes from the fact that there are two sites with the same distance in 1D. The result of this calculation is shown in Fig. 3(c) in the main text.